The sum of two teachers is 98. One of the teachers is 2 years older than the other. How old are both of the teachers?
Answer by addingup(3677)

Let the younger teacher be x.
Then the older teacher will be x+2
----------------------------------------
x+x+2= 98
2x+2= 98
2x= 96
x= 43 The younger one is 43 and the other:
43+2= 45
J